CURRICULUM VITAE

Name: Kari-Pekka ENQVIST

Date and place of birth: 16th of February, 1954, Lahti, Finland

Current position: Senior Research Fellow, Academy of Finland 1.8.1996-31.7.2001; Senior Research Associate, Physics Department, University of Helsinki (on a leave of absence); Cosmology Project Leader, Helsinki Institute of Physics 1998-1999; Theory Programme Leader, Helsinki Institute of Physics 1999-

Education and training:

• graduated from the University of Helsinki, Finland, April 1980, majoring in physics with minors in theoretical physics, mathematics, astronomy and philosophy (additional degrees in psychology and social psychology)

• Licentiate of Philosophy (a postgraduate intermediate degree), May 1981, at the University of Helsinki

• Ph.D. in Elementary Particle Physics at the University of Helsinki, May 1983 (supervisor prof. Masud Chaichian, University of Helsinki).

Previous professional appointments:

• 1978–1983 Research assistant at the Department of High Energy Physics and the Research Institute for Theoretical Physics, University of Helsinki, Finland.

• 1984–1986 Research fellow at CERN, TH division, Geneva, Switzerland (24 months).

• 1986–1987 Research fellow at the Physics Department, University of Wisconsin, Madison, USA (12 months).

• 1987–1990 Senior research fellow at the Research Institute for Theoretical Physics, University of Helsinki (22 months).

• 1988–1989 Acting associate professor at the Department of Theoretical Physics, University of Helsinki (12 months).

• 1990–1994 Assistant professor, Nordita, Copenhagen, Denmark (49 months).

• 1995-1996 Acting professor, Department of Physics, University of Helsinki (17 months).

• 1996- Senior research fellow, Academy of Finland.

• Docent of Theoretical Physics at the University of Helsinki, December 1984; Research Associate, Dept. of Theoretical Physics, University of Helsinki, 1985-1995; Senior Research Associate, Theory Division, Physics Department, University of Helsinki (on a leave of absence), 1996- • Co-Investigator in the Low Frequency Instrument Consortium of the ESA Planck Surveyor Mission 1998-

Research awards, honours, stipends: Valtion Tiedonjulkistamispalkinto 1995; Lauri Jäntti foundation award 1997; The Magnus Ehrnrooth Foundation Physics Award 1997; Tieto-Finlandia 1999

Other academic and professional activities:

• 146 scientific publications (in English); total number of citations in the SLAC QSPIRES database: 2909. 5 most cited papers in the list of publications are: 42 (cited 248 times), 38 (cited 198 times) 45, (cited 170 times), 94 (cited 125 times) and 18 (cited 121 times).

• Organiser of: Nordita Summer School (Porvoo, June 1989); Nordita Workshop on Electroweak Matter (Copenhagen, January 1991); Nordita Program of Neutrino Physics (Copenhagen, May-June 1992); Nordic-Baltic Summer School (Lohusalu, Estonia, July 1992); Workshop on the Electroweak Phase Transition (Helsinki 1994); Summer School and Mini-Workshop in Astroparticle Physic (Denmark 1995); Neutrino 96 (Helsinki 1996) EPS-99 (Tampere 1999)

• Member of the Nordic Network ”The Standard Model in Extreme Environments”. Coordinator of the EU Training and Mobility programme “A Critical Investigation of Electroweak Baryogenesis Models”.

• Referee for Phys. Rev. D, Phys. Rev. Lett., Phys. Lett. B, Nucl. Phys. B, Z. für Physik C, Physica Scripta, Mod. Phys. Lett., Europhysics Letters

• In the International Advisory Committee of: 2nd Tallinn Symposium on Neutrino Physics (Estonia 1993); Trends in Astroparticle Physics (Sweden 1994); Nordic Symposium ’Cosmology: from COBE to galaxy formation’ (Denmark 1997); Nordic Winter School on Particle Physics and Cosmology (Norway, January 2-8 1999); Advisor to the Research Council of Norway; Advisor to the Academy of Finland.

• recent conference attendance: Moriond (Les Arcs) March 1990; Neutrino 90 (CERN) June 1990; 2nd Workshop on Finite Temperature Field Theories (Tsukuba) July 1990; 25th Int’l HEP Conference HEP (Singapore) August 1990; Invited talk at the Tallinn Symposium on Neutrino Physics (Lohusalu, Estonia) October 1990; Lepton-Photon 91-HEP Conference (Geneva), July 1991; Invited talk at the Int. Workshop on Electroweak Physics Beyond the Standard Model, (Valencia), October 1991; Invited talk at the Nordic Meeting on Particle Physics (Spåtind, Norway), January 1992; Invited talk at the XXVI Int’l HEP Conference (Dallas), August 1992; Invited talk at the Kazimierz meeting (Poland), May 1993; Invited talk at Electroweak Phase Transitions (Sintra, Portugal) March 1994; Neutrino 94 (Eilat, Israel), June 1994; Invited talk at the XXVII Int’l HEP Conference (Glasgow), July 1994; Trends in Astroparticle Physics (Stockholm) September 1994; TAN network meeting (Annecy) November 1994; 4th International Workshop on Theoretical and Phenomenological Aspects of Underground Physics (Toledo) September 1995, Uppsala/Nordita Workshop in Astroparticle Physics, April 1996; Invited talk at Strong and Eleectroweak Matter ’97, (Eger, Hungary), May 1997; Invited talk at Cosmo 97-conference (Ambleside, Great Britain), September 1997; Invited talk at Nordic Symposium ’Cosmology: from COBE to galaxy formation’ (Copenhagen), December 1997; Invited talk at PAS-COS 98 Conference (Boston) March 1998; Invited talk at Hot Non-Perturbative Particle Physics Network meeting (Copenhagen) April 1998; Invited talk at the Planck Miniworkshop (Helsinki University) May 1998; Invited talk at CAPP-98 Conference (CERN), June 1998; Talk at SUSY98 Conference (Oxford), July 1998; Invited talk at DARK98 Conference (Heidelberg), July 1998; Invited talk at Strong and Electroweak Matter conference (Copenhagen) December 1998; Invited talk at COSMO99 Conference (Trieste), September 1999.

Other activities

• Vice president of the Finnish Sceptical Society SKEPSIS 1996-1998; board member of the Finnish Amateur Astronomical Society URSA 1999-; subeditor of the Finnish Physical Society Journal Arkhimedes; member of the advisory board of the University of Helsinki journal Yliopisto; member of the Finnish Unesco Committee (Section of Natural Sciences) 1999-

• about 90 popular articles (in finnish), including the books Tyhjästä syntynyt (WSOY 1994, with J. Maalampi; short listed for the annual Tieto-Finlandia Award); Näkymätön todellisuus (WSOY 1996; winner of the Lauri Jäntti foundation award; short listed for the annual Tieto-Finlandia Award) and Olemisen porteilla (WSOY 1998) (winner of the annual Tieto-Finlandia Award).